A boat travels 30 mi upstream and 30 mi downstream. The total time for both parts of the trip is 8 hr. The speed of the boat in still water is 8 mph. Find the speed of the current.
Answer by rwm(914)

a+b=8
(8+c)*a=30
(8-c)*b=30
a=3 hrs with the current
b=5 hrs against the current
current c=2
check
10*3=30
6*5=30
5+3=8
ok